OpenStudy (anonymous):

Please help me with this problem. @jim_thompson5910 24/36 = 16/x ?

jimthompson5910 (jim_thompson5910):

first step is to cross multiply

jimthompson5910 (jim_thompson5910):

what do you get

OpenStudy (anonymous):

Is the equation for cross multiplication above correct?

jimthompson5910 (jim_thompson5910):

yes you have the right setup, you now need to solve

OpenStudy (anonymous):

x= 24

jimthompson5910 (jim_thompson5910):

yes

OpenStudy (anonymous):

X=24 If you don't understand, go to google and search : "solve for x calculator" and it will explain after you enter the equation! :)

OpenStudy (anonymous):

Thanks .

OpenStudy (anonymous):

Yep! :)

OpenStudy (anonymous):

@jim_thompson5910 Now what? This is where I get stuck.

jimthompson5910 (jim_thompson5910):

x = 24 is the length of BD

jimthompson5910 (jim_thompson5910):

that ratio you set up above corresponds to the lengths of the triangles

OpenStudy (anonymous):

None of the options are 24 :(

jimthompson5910 (jim_thompson5910):

ok so we can't use BD

jimthompson5910 (jim_thompson5910):

try to find the length of BF

jimthompson5910 (jim_thompson5910):

hint: BF = DE (because BDEF is a parallelogram)

OpenStudy (anonymous):

36?

jimthompson5910 (jim_thompson5910):

focus on triangles ADE and EFC

jimthompson5910 (jim_thompson5910):

use them to find DE

OpenStudy (anonymous):

x/16= 36/27 ?

OpenStudy (anonymous):

Are you trying to solve for x?

jimthompson5910 (jim_thompson5910):



jimthompson5910 (jim_thompson5910):

 so we have 24/36 = x/27